OpenStudy (anonymous):

find x (x-5)/(3-x)<0 and no the answer isn't x<5

OpenStudy (amistre64):

what else aint it then?

OpenStudy (anonymous):

you see if you plug 4 in doesnt give you a number less than 0

OpenStudy (amistre64):

\[\frac{x-5}{3-x}<0\] means its negative when x = 5 we have 0 when 3 = 3 we have a bad term but useful nonetheless <................ 3 .....................5...............> - + + - - + ------------------------------ + - + between 3 and 5 it looks like

OpenStudy (amistre64):

in interval wed write: (3,5)

OpenStudy (anonymous):

so the answer is 3 and 5?

OpenStudy (amistre64):

with an inequality you get a "range" of answers

OpenStudy (amistre64):

it is all the values between 3 and 5

OpenStudy (anonymous):

but 4 doent work!!! :(

OpenStudy (amistre64):

show me why

OpenStudy (amistre64):

i see it :)

OpenStudy (anonymous):

do u get what im saying?

OpenStudy (amistre64):

yep, i missed something along the way; probably a negative switch http://www.wolframalpha.com/input/?i=%28x-5%29%2F%283-x%29+%3C0

OpenStudy (anonymous):

howd they get x>5?

OpenStudy (amistre64):

x-5 / 3-x < 0 , the criticals are 3 and 5 0-5 / 3-0 < 0 -/+ = - 6-5 / 3-6 < 0 +/- = - .................................. so all point below 3 and above 5 are in the zone; but thats just by logic, let me type up the algebra to be sure